The Palmer Acres Inn is trying to determine its break-even point during its off-peak season. The inn has 50 rooms that it rents at $ 70  a night. Operating costs are as follows.

Salaries $ 5,800 per month
Utilities $ 1,400 per month
Depreciation $ 1,100 per month
Maintenance $ 3,740 per month
Maid service $ 14 per room
Other costs $ 28 per room

Determine the inn’s break-even point in number of rented rooms per month.

Break-even point

enter the break-even point in number of rented rooms per month

rooms

eTextbook and Media

  

  

Determine the inn’s break-even point in dollars.

Break-even point

$enter the break-even point in dollars
Answer #1

Total variable cost=Maid service+Other costs

=(14+28)=$42 per room

Total fixed cost=Salaries+Utilities+Depreciation+Maintenance

=(5800+1400+1100+3740)=$12040

Contribution margin=Revenue-Variable cost  

=(70-42)=$28 per room

Breakeven=Fixed expenses/Contribution margin

=(12040/28)=430 rooms

=(430*70)=$30100